OpenStudy (anonymous):

Simplify the expression. √27x^5 + x√3 + 9√3x If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (pfenn1):

Is this your expressioon?\[\sqrt{27} x^5 + x \sqrt{3} + 9\sqrt{3x}\]

OpenStudy (anonymous):

Yes.

OpenStudy (pfenn1):

Are you sure it isn't\[\sqrt{27} x^5 + x \sqrt{3} + 9\sqrt{3}x\]

OpenStudy (pfenn1):

\[\sqrt{27x^5} + x \sqrt{3} + 9\sqrt{3x}\]

OpenStudy (anonymous):

Oh, sorry. You're right.

OpenStudy (pfenn1):

Which is it? The first, second or third expression?

OpenStudy (anonymous):

The third one.

OpenStudy (pfenn1):

Okay. So our expression is \[\sqrt{27x^5} + x \sqrt{3} + 9\sqrt{3x}\]Can you simplify \[\sqrt{27x^5}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

It's 3x^2√3x....

OpenStudy (pfenn1):

sure, so what can you to simplify the whole initial expression? Sorry, I have to run as dinner is on the table.

OpenStudy (pfenn1):

\[3x^2\sqrt{3x}+x \sqrt{3}+9\sqrt{3x}\]